# TME-Box: Scalable In-Process Isolation through Intel TME-MK Memory Encryption

### v1

@@ V-B Memory Allocator
Furthermore, the memory allocator is responsible for initializing and managing heap memory. Therefore, the memory allocator maps the heap for every sandbox (with their designated keyID) and aliases it to the same physical memory region. This procedure enables the memory allocator to apply the scalable isolation and flexible relocation of data in memory. Precisely, the memory allocator can use fine-grained isolation, ranging from individual cache lines to a large number of pages, depending on the runtime constraints. On pages that contain data from multiple sandboxes, the allocator needs to align memory chunks to cache line size. Here, our design enables the relocation of data, as individual memory chunks can be immediately reallocated. Note that data from different sandboxes is never co-located within a single cache line.
@@ VI-A Systematic Analysis
The TME-MK encryption engine provides integrity protection for DRAM memory. Once a memory location (i.e., cache line) is initialized with a specific keyID, it is protected by the MAC, which preserves integrity at the cache line level. During memory reads, the memory controller requests the DRAM data and its associated MAC, verifying the memory access. Successfully verified memory accesses can result in the caching of the data, with the corresponding cache line marked with the keyID used for the access. Any attempt to access memory with an incorrect keyID is detected by the MAC verification within cryptographic bounds. Writing to memory with an incorrect keyID corrupts the associated MAC. Although this will not immediately result in the detection of the violation, this corruption is also not security critical since no secret information can be extracted. Subsequent memory reads from this corrupted memory location will trigger a MAC authentication, thus detecting the corruption of the sandbox’s memory. In such cases, the trusted runtime handles the exception, preventing the attacker from exploiting the system.
@@
Moreover, software sandboxing typically also ensures that untrusted sandboxed code within the process has only restricted access to system calls. These system calls are then executed by a trusted runtime, which can be invoked by the sandbox through a defined interface. This includes critical system calls, such as mprotect and mmap, used for managing memory resources. Another attack path for memory corruption consists of gaining an arbitrary read and write primitive followed by remote code execution (RCE), i.e., by invoking the exec system call. This is followed by a privilege escalation attack to compromise the entire software system. To mitigate this security threat, the usage of system calls is only allowed for the trusted runtime, i.e., the sandbox has to invoke system calls through the trusted runtime. Also, the no-execution policy on writable memory (i.e., Write-XOR-Execute) prevents code injection attacks [6]. In addition, orthogonal measures such as syscall filtering [67] can be employed to further restrict access to system calls.
@@
When performing data relocation, the memory location is initialized with a new keyID, forcing a writeback for the data in memory that also updates the MAC. This procedure ensures that access for sandboxes with revoked keyIDs becomes invalid. As a result, any subsequent attempts to access memory locations with the revocated keyID trigger MAC verification errors. Note that the processor ensures cache and TLB coherency, maintaining consistency for the cached data and keyIDs located in the PTEs across all CPU cores.
@@ VII-A Performance Evaluation
Additionally, we perform an evaluation of the NGINX web server (version 1.26.0) using an experimental setup comparable to prior work [84, 81, 58]. This experiment uses ApacheBench (ab) to generate requests to receive files of different sizes from the NGINX web server. We compile NGINX with our TME-Box compiler toolchain and run a single NGINX worker pinned to an isolated CPU core. We use ApacheBench to perform and benchmark 2,000,000 requests from one client with increasing file sizes covering , , , , and . Figure 9 shows the throughput of NGINX with TME-Box in both modes (gs and r15) for code and data isolation normalized to the performance of an unmodified NGINX version. Here, the decrease in throughput for NGINX ranges from 8.3 % to 1.8 % for TME-Box (gs) and from 8.7 % to 1.9 % for TME-Box (r15). We confirm the observations of prior work [84, 81] that the measured throughput decreases stronger for smaller file sizes than for larger files, i.e., the overhead declines as the file size increases.
@@ VIII-A Related Work
However, the isolation of individual software components with heavy-weight protection mechanisms, such as process isolation, has an impact on system performance and start-up times [3]. Thus, process isolation is not well suited for the isolation of individual software components of applications. As a result, some software systems, like cloud worker architectures [17], require more flexible and efficient in-process isolation that allows scalable memory isolation for different sandboxes, as provided by our mechanism.
@@
Protection keys for user space (PKU) is an approach for memory isolation that relies on additional metadata stored in page table entries (PTE). The Intel memory protection keys (MPK) [61] allow to dynamically assign a 4-bit protection key to a specific page. For each memory operation, logical integrity checks are performed that compare the page’s protection key against the active access policy located and controlled by the user space protection key register (PKRU). The current access permissions (i.e., read and write) of each individual protection key can be changed from user space via the wrpkru instruction. In this way, MPK can be used for data protection, e.g., for dynamically locking-away of sensitive data in memory like cryptographic key material.
Based on MPK, several academic designs propose software sandboxing [32, 81, 67] that facilitates in-process isolation. However, PKU-based sandboxing can have performance implications in certain scenarios that require frequent sandbox transitions since the wrpkru instruction (including necessary fences) can take more than 100 cycles [81, 68]. Similarly, ARMlock [88] enables in-process isolation using the, meanwhile deprecated, ARM memory domains (that provide a 4-bit domain ID for isolation). However, ARM’s policy changes need to be performed by the kernel as the policy register is not accessible from user space, incurring additional runtime overhead for policy changes.
Intel MPK and ARM memory domains only allow for 16 distinct domains, which can be restrictive for certain software architectures that require more in-process domains. Also, both ARM memory domains and Intel MPK are constrained to the granularity of a page and cannot achieve sub-page granular isolation. In contrast to protection keys, TME-Box allows the isolation of more in-process domains and provides fine-grained memory isolation on the level of individual cache lines.
@@
Memory tagging technologies associate metadata with data located in memory at a certain granularity, thereby enabling the enforcement of different security policies that restrict access to memory resources [41]. Tagged architectures such as the ARM memory tagging extension (MTE) [71] and SPARC application data integrity (ADI) [4] hardware features have been efficiently integrated into commercial processors. ARM MTE is typically used to provide probabilistic memory safety for memory sanitization (e.g., MemTagSanatizer) or to help in establishing runtime security for production software. However, ARM MTE and SPARC ADI are often criticized for their small tag size of 4-bit (i.e., 16 distinct memory tags), resulting in a relatively low detection probability of just 93.75 % [72].
Furthermore, SFITAG [70] combines software instrumentation with ARM MTE to isolate kernel extensions. However, their design only supports a maximum of 14 isolated memory regions, as two memory tags are reserved. In contrast, our TME-Box design supports a larger number of concurrent sandboxes (i.e., TME-MK is specified for up to 15-bit keyIDs that address 32K encryption keys [34]). Also, HAKC [52] combines ARM pointer authentication (PAuth) [79] with ARM MTE for kernel compartmentalization. While HAKC’s design expands the possible number of compartments from 4-bit MTE, their security is bound to the cryptographic MAC of the PAuth feature, which depends on the size of the virtual address space. More specifically, ARM allows the configuration of virtual address sizes between 32 and 52 bits (with bit 55 reserved). This means that a large virtual address space, as used by servers, can lead to a relatively small MAC, e.g., a 52-bit virtual address space results in a MAC size of 11 bits, or even just 3 bits if ARM’s top-byte ignore (TBI) is enabled [79].
@@
Intel TME-MK can be used for runtime security to mitigate various attack vectors [66, 65, 60, 13]. For example, IntegriTag [66] uses Intel TME-MK for probabilistic heap memory safety. IntegriTag implicitly encodes a pseudorandomly selected keyID into the virtual address and assigns the keyID to the corresponding memory location through aliasing. This method restricts access to heap objects, as memory accesses are only granted by pointers that incorporate the correct keyID. This way, TME-MK can be used in the same way as memory tagging, like ARM MTE and SPARC ADI, providing similar security. Nevertheless, adversaries may still exploit vulnerabilities to leak or guess the keyID, and to harvest or forge valid pointers, thereby circumventing the security measure. In contrast, TME-Box enforces the use of the sandbox’s keyID for all memory operations, enabling scalable in-process isolation. Moreover, EC-CFI [60] uses Intel TME-MK to thwart fault-induced control-flow hijacking attacks. EC-CFI encrypts code at the function level, enforcing decryption with different keys for individual functions. These encryption keys are switched when entering and exiting a function, thereby applying control-flow protection in the presence of fault attacks. EC-CFI helps to prevent fault-induced control flow hijacking attacks, while our work focuses on software sandboxing.
In addition, Intel trust domain extensions (TDX) [13] enable confidential computing with heavy-weight virtual machine (VM) isolation. TDX uses TME-MK for the encryption of VMs and containers [36], and the protection against physical attacks [35, 31]. In contrast, TME-Box allows for lightweight and efficient in-process sandboxing without relying on the use of heavy-weight virtualization mechanisms.
@@ VIII-B Possible Extensions
The TME-Box design can be adopted for the isolation of just-in-time (JIT) compiled code since in-process isolation is highly demanded in this context, as seen in the V8 sandbox [29, 28]. To achieve this, the JIT compiler must be TME-Box aware to effectively enable our isolation and memory management, i.e., the JIT compiler must guarantee that the required instrumentation is performed similarly to our compiler extension. Furthermore, the JIT compiler is responsible for only emitting safe instructions. Moreover, the JIT compiler has to adhere to the Write-XOR-Execute policy, which we assume is enabled for our design. For instance, the Google Native Client (NaCl) [86, 69] sandbox has been extended to the JIT compiler of the V8 engine [5]. We leave it as future work to port TME-Box to JIT compilers.
@@ IX Conclusion
TME-Box extends the application of TME-MK’s integrity enforcement to cryptographically isolate the memory of sandboxes within a single process (i.e., in-process isolation). That is, TME-Box uses compiler instrumentation to enforce that the sandboxes use their designated encryption keys for memory interactions, thus detecting unauthorized memory accesses. Repurposing TME-MK’s runtime encryption for hardware-assisted sandboxing provides unique isolation properties: TME-Box is scalable, enabling isolation granularities from individual cache lines up to full pages. Moreover, our design allows the isolation for up to 32K concurrent sandboxes and offers flexible memory management, i.e., data relocation in memory.
@@ Acknowledgment
We thank the anonymous reviewers for their valuable feedback that improved this work. This project has received funding from the Austrian Research Promotion Agency (FFG) via the AWARE project (FFG grant number 891092) and the SEIZE project (FFG grant number 888087). Additional funding was provided by a generous gift from Intel. Any opinions, findings, conclusions, or recommendations expressed in this paper are those of the authors and do not necessarily reflect the views of the funding parties.

### v2

@@ V-B Memory Allocator
Furthermore, the memory allocator is responsible for initializing and managing heap memory. Therefore, the memory allocator maps the heap for every sandbox (with their designated keyID) and aliases it to the same physical memory region. Moreover, the memory allocator then performs the initialization of the corresponding heap memory on allocation. This procedure enables the memory allocator to apply the scalable isolation and flexible relocation of data in memory. Precisely, the memory allocator can use fine-grained isolation, ranging from individual cache lines to a large number of pages, depending on the runtime constraints. On pages that contain data from multiple sandboxes, the allocator needs to align memory chunks to cache line size. Here, our design enables the relocation of data, as individual memory chunks can be immediately reallocated. Note that data from different sandboxes is never co-located within a single cache line.
@@ VI-A Systematic Analysis
The TME-MK encryption engine provides integrity protection for DRAM memory. Once a memory location (i.e., cache line) is initialized with a specific keyID, it is protected by the MAC, which preserves integrity at the cache line level. During memory reads, the memory controller requests the DRAM data and its associated MAC, verifying the memory access.
Successfully verified memory accesses can result in the caching of the data, with the corresponding cache line marked with the keyID used for the access. Any attempt to access memory with an incorrect keyID is detected by the MAC verification within cryptographic bounds. Specifically, cache lines are tagged with the physical address, and the keyID is a part of that. Thus, accessing cache lines with a different keyID results in a cache miss and leads to a DRAM access. This DRAM access is then authenticated by the encryption engine, which detects the usage of incorrect keyIDs.
Writing to memory with an incorrect keyID corrupts the associated MAC. Although this will not immediately result in the detection of the violation, this corruption is also not security critical since no secret information can be extracted. Subsequent memory reads from this corrupted memory location will trigger a MAC authentication, thus detecting the corruption of the sandbox’s memory. In such cases, the trusted runtime handles the exception, preventing the attacker from exploiting the system.
@@
Moreover, software sandboxing typically also ensures that untrusted sandboxed code within the process has only restricted access to system calls. These system calls are then executed by a trusted runtime, which can be invoked by the sandbox through a defined interface. This includes critical system calls, such as mprotect and mmap, used for managing memory resources. Another attack path for memory corruption consists of gaining an arbitrary read and write primitive followed by remote code execution (RCE), e.g., by invoking the exec system call. This is followed by a privilege escalation attack to compromise the entire software system [91, 49, 47]. To mitigate this security threat, the usage of system calls is only allowed for the trusted runtime, i.e., the sandbox has to invoke system calls through the trusted runtime. Also, the no-execution policy on writable memory (i.e., Write-XOR-Execute) prevents code injection attacks [6]. In addition, orthogonal measures such as syscall filtering [71] can be employed to further restrict access to system calls.
@@
When performing data relocation, the memory location is initialized with a new keyID, forcing a writeback for the data in memory that also updates the MAC. This procedure ensures that access for sandboxes with revoked keyIDs becomes invalid. As a result, any subsequent attempts to access memory locations with the revoked keyID trigger MAC verification errors. Note that the processor ensures cache and TLB coherency, maintaining consistency for the cached data and keyIDs located in the PTEs across all CPU cores.
@@ VII-A Performance Evaluation
Additionally, we perform an evaluation of the NGINX web server (version 1.26.0) using an experimental setup comparable to prior work [87, 84, 58]. This experiment uses ApacheBench (ab) to generate requests to receive files of different sizes from the NGINX web server. We compile NGINX with our TME-Box compiler toolchain and run a single NGINX worker pinned to an isolated CPU core. We use ApacheBench to perform and benchmark 2,000,000 requests from one client with increasing file sizes covering , , , , and . Figure 9 shows the throughput of NGINX with TME-Box in both modes (gs and r15) for code and data isolation normalized to the performance of an unmodified NGINX version. Here, the decrease in throughput for NGINX ranges from 8.3 % to 1.8 % for TME-Box (gs) and from 8.7 % to 1.9 % for TME-Box (r15). We confirm the observations of prior work [87, 84] that the measured throughput decreases stronger for smaller files than for larger files, i.e., the overhead declines as the file size increases.
@@ VIII-A Related Work
However, the isolation of individual software components with heavy-weight protection mechanisms, such as process isolation, has an impact on system performance and start-up times [3, 66]. Thus, process isolation is not well suited for the isolation of individual software components of applications. As a result, some software systems, like cloud worker architectures [17], require more flexible and efficient in-process isolation that allows scalable memory isolation for different sandboxes, as provided by our mechanism.
@@
Protection keys for user space (PKU) is an approach for memory isolation that relies on additional metadata stored in page table entries (PTE). The Intel memory protection keys (MPK) [61] allow to dynamically assign a 4-bit protection key to a specific page. For each memory operation, logical integrity checks are performed that compare the page’s protection key against the active access policy located and controlled by the user space protection key register (PKRU). The current access permissions (i.e., read and write) of each individual protection key can be changed from user space via the wrpkru instruction. In this way, MPK can be used for data protection, e.g., for dynamically locking-away sensitive data in memory like cryptographic key material.
Based on MPK, several academic designs propose software sandboxing [31, 84, 62, 72, 71, 8] that facilitates in-process isolation. However, PKU-based sandboxing can have performance implications in certain scenarios that require frequent sandbox transitions since the wrpkru instruction (including necessary fences) can take more than 100 cycles [84, 72]. Similarly, ARMlock [93] enables in-process isolation using the, meanwhile deprecated, ARM memory domains (that provide a 4-bit domain ID for isolation). However, ARM’s policy changes need to be performed by the kernel as the policy register is not accessible from user space, incurring additional runtime overhead for policy changes.
Intel MPK and ARM memory domains only allow for 16 distinct domains, which can be restrictive for certain software architectures that require more in-process domains. Also, both Intel MPK and ARM memory domains are constrained to the granularity of a page and cannot achieve sub-page granular isolation. In contrast to protection keys, TME-Box allows the isolation of more in-process domains and provides fine-grained memory isolation on the level of individual cache lines.
@@
Memory tagging technologies associate metadata with data located in memory at a certain granularity, thereby enabling the enforcement of different security policies that restrict access to memory resources [38].
Tagged architectures such as the ARM memory tagging extension (MTE) [75] and SPARC application data integrity (ADI) [4] hardware features have been efficiently integrated into commercial processors. ARM MTE is typically used to provide probabilistic memory safety for memory sanitization (e.g., MemTagSanatizer) or to help establish runtime security for production software. However, ARM MTE and SPARC ADI are often criticized for their small tag size of 4-bit (i.e., 16 distinct memory tags), resulting in a relatively low detection probability of just 93.75 % [76].
Furthermore, SFITAG [74] combines software instrumentation with ARM MTE to isolate kernel extensions. However, their design only supports a maximum of 14 isolated memory regions, as two memory tags are reserved. In contrast, our TME-Box design supports a larger number of concurrent sandboxes (i.e., TME-MK is specified for up to 15-bit keyIDs that address 32K encryption keys [32]).
Also, HAKC [52] combines ARM pointer authentication (PAuth) [64] with ARM MTE for kernel compartmentalization. While HAKC’s design expands the possible number of compartments from 4-bit MTE, their security is bound to the cryptographic MAC of the PAuth feature, which depends on the size of the virtual address space. More specifically, ARM allows the configuration of virtual address sizes between 32 and 52 bits (with bit 55 reserved). This means that a large virtual address space, as used by servers, can lead to a relatively small MAC, e.g., a 52-bit virtual address space results in a MAC size of 11 bits, or even just 3 bits if ARM’s top-byte ignore (TBI) is enabled [64].
@@
Intel TME-MK can be used for runtime security to mitigate various attack vectors [70, 69, 60, 15]. For example, IntegriTag [70] uses Intel TME-MK for probabilistic heap memory safety. IntegriTag uses different security policies (e.g., pseudorandom keyIDs) that select and implicitly encode a keyID into the pointer’s virtual address and assign the keyID to the corresponding memory location through aliasing. This method restricts access to heap objects, as memory accesses are only granted by pointers that incorporate the correct keyID. In this way, TME-MK can be used in the same way as memory tagging, like ARM MTE and SPARC ADI, providing similar security. Nevertheless, adversaries may still exploit vulnerabilities to leak or guess the keyID, and to harvest or forge valid pointers, thereby circumventing the security measure. In contrast, TME-Box enforces the use of the sandbox’s keyID for all memory operations, enabling scalable in-process isolation. Moreover, EC-CFI [60] uses Intel TME-MK to thwart fault-induced control-flow hijacking attacks. EC-CFI encrypts code at the function level, enforcing decryption with different keys for individual functions. These encryption keys are switched when entering and exiting a function, thereby applying control-flow protection in the presence of fault attacks. EC-CFI helps to prevent fault-induced control flow hijacking attacks, while our work focuses on software sandboxing.
In addition, Intel trust domain extensions (TDX) [15, 33] enable confidential computing with heavy-weight virtual machine (VM) isolation. TDX uses TME-MK for the encryption of VMs and containers [34], and the protection against physical attacks [33, 30]. In contrast, TME-Box allows for lightweight and efficient in-process sandboxing without relying on the use of heavy-weight virtualization mechanisms.
@@ VIII-B Possible Extensions
The TME-Box design can be adopted for the isolation of just-in-time (JIT) compiled code since in-process isolation is highly demanded in this context, as seen in the V8 sandbox [29, 27, 68]. To achieve this, the JIT compiler must be TME-Box aware to effectively enable our isolation and memory management, i.e., the JIT compiler must guarantee that the required instrumentation is performed similarly to our compiler extension. Moreover, the JIT compiler has to adhere to the Write-XOR-Execute policy [92, 68], which we assume is enabled for our design. For instance, the Google Native Client (NaCl) [89, 73] sandbox has been extended to the JIT compiler of the V8 engine [5]. We leave it as future work to port TME-Box to JIT compilers.
@@ IX Conclusion
TME-Box extends the application of TME-MK’s integrity enforcement to cryptographically isolate the memory of sandboxes within a single process (i.e., in-process isolation). That is, TME-Box uses compiler instrumentation to enforce that the sandboxes use their designated encryption keys for memory interactions, thus detecting unauthorized memory accesses. Repurposing TME-MK’s runtime encryption for hardware-assisted sandboxing provides unique isolation properties: TME-Box is scalable, enabling isolation granularities from individual cache lines up to full pages. Moreover, our design allows the isolation for up to 32K concurrent sandboxes and offers flexible memory management for the allocator, i.e., data relocation in memory.
@@ Acknowledgment
We thank Andreas Kogler and the anonymous reviewers for their valuable feedback that improved this work. This project has received funding from the Austrian Research Promotion Agency (FFG) via the AWARE project (FFG grant number 891092) and the SEIZE project (FFG grant number 888087). Additional funding was provided by a generous gift from Intel. Any opinions, findings, and conclusions or recommendations expressed in this paper are those of the authors and do not necessarily reflect the views of the funding parties.